BI shutdown follow-up meeting

Europe/Zurich
CERN

CERN

Ray - IEFC highlights

Ray shows schedules for 2021 in the injectors and LIU commissioning

Christos - MD in SPS will define readiness dates for equipment

Manfred - no requests for MD but this will happen.

 

PS complex (William)

Needs input on LS2 feedback and plans for 2021/22 with a talk planned for next week. Only 3 answers so far for interventions in YETS 21/22 and 2 for feedback.

Need inputs, in particular for vacuum interventions.

 

SPS (Morad)

Also needs input for LS2 and for YETS interventions.

Thibaut comments that we will have more input during the year.

TT10 fibres planned for shutdown - this is being discussed with EN, but not formally done.

 

LHC (Gerhard)

Most sectors cooling or cold.

BLMs following planning. BTV, BRAN maintenance done

BCTs - still some parts for analogue system in galeries (Tom Levens)

Manfred - will install in 6 arc the prototype BPMs in the alcove later in the year.

BPM activities (Michal). Some were waiting for bakeout and others were outliers.

Needs list of activities for this year, before and after beam tests.

 

Ana M - month-by-month interventions

Schedules are no longer updated, so info from IMPACT and other sources.

AD to be added.

Perhaps stop doing these monthly schedules

 

PSB commissioning update (Michele)

Michele shows status of the instrument commissioning.

William comments that the noise on the old BWS in rings 1,2 has been fixed.

Diogo - noisy (ring) BPMs in specific regions of the cycle. Is this the distributor? Michele - still under investigation, but could be the same problem as for the BCTs.

 

PS readiness for beam (Ana G)

BSG for NToF needed, but all others tested OK.

BGIV HV issue with new box being prepared and cooling pipes to be installed.

BLMs (ring and TT2) some changes on SW and FW ongoing. Should be deployed soon.

BWS installation still ongoing, with cabling to be finished along with acquisition electronics. Resources are committed to PSB for now. Some problem with the acquisition in the PSB, still improving FW and SW. Still 1 month for PS, and should be a copy-paste of the PSB.

First beam 1st March.

Will

 

SPS Commissioning (Tom)

Tom is the new BI-SPS Coordinator

Most work on RF and magnets for now.

Christos - Issues this week with a cable missing (removed in error) for the interlock loop. Not sure why this happened. This prevents beams into the NA, so critical! Alternatives also being investigated.

TDC2 access ongoing, ALARA2 with ~1mSv collective. Some cables also broken. Not sure yet of causes.

There are minutes attached to this event. Show them.